OpenStudy (anonymous):

Solve the following system of equations and show all of your work. 2x + 5y = −3 −15x − 5y = −10

OpenStudy (anonymous):

need help please : (

Directrix (directrix):

2x + 5y = −3 −15x − 5y = −10 Let's use addition on this. -13x +0y = -13 -13x = - 13 x = 1

Directrix (directrix):

To get y, pick either equation and substitute 1 for x and then get y. 2x + 5y = −3 2(1) + 5y = -3 2 + 5y = -3 5y = -3 -2 5y = -5 y = -1 Solution (1, -1)

OpenStudy (anonymous):

this is a system by elimination because in the 1rt question I have to choose

OpenStudy (anonymous):

Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method.

Directrix (directrix):

Is this a new question or does it go with the first problem?

OpenStudy (anonymous):

can you help me whit this one because I chose substitution and now I'm not sure

OpenStudy (anonymous):

is same question

OpenStudy (anonymous):

it goes with the first problem

OpenStudy (anonymous):

are u confuse

OpenStudy (anonymous):

???

Directrix (directrix):

You can use any of those methods to solve the system. When you used substitution, did you get the correct answer?

Directrix (directrix):

Do you want me to explain in sentences why I chose the elimination technique?

OpenStudy (anonymous):

yes please but I dont have to chose one of this method to graph

Directrix (directrix):

I looked at the two equations 2x + 5y = −3 and −15x − 5y = −10 and saw that the y terms were additive inverses and would add to 0 and eliminate the y-variable. With the y-variable gone, I would be left with one equation in one unknown which I could solve. That is -13x +0y = -13 which has an answer of x = 1.

OpenStudy (anonymous):

oh ok get you now make sense now =) thank you

Directrix (directrix):

Then, I had to find y by substituting 1 for x in one of the two given equations. I chose the first one. 2x + 5y = −3 and got -1 for y. See below. 2(1) + 5y = -3 2 + 5y = -3 5y = -3 -2 5y = -5 y = -1 So, the solution for the system is (1,-1).

Directrix (directrix):

What do we do next?